OpenStudy (anonymous):

anyone can explain how this? ( (1/(1-2sin²θ)))-1 / ((2sinθ)/(1-2sinθ)) <-----?? = sinθ become this? ((1-1+2sin²θ)/(2sinθ))=sinθ

OpenStudy (zehanz):

Is this the expression you've got:\[\frac{ 1 }{ 1-2\sin^2 \theta }-\frac{ 1 }{ \frac{ 2\sin \theta }{ 1-2\sin \theta } }=\frac{ 1 }{ 1-2\sin^2 \theta }-\frac{ 1-2\sin \theta }{ 2\sin \theta }\]

OpenStudy (anonymous):